OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. B(6, 2), m = -1/2

OpenStudy (anonymous):

since you have a point, and the slope of a line, you can use point-slope formula, which is: \[(y-y _{1})=m(x-x _{1})\] your x1 and y1 are the given point, and the m is the slope simply insert your numbers \[y-2=(-1/2)(x-6)\] you may want your answer in y=mx+b form so all you do is solve for y distribute the (-1/2) \[y-2=(-1/2x)+3\] then just add 2 to both sides y=(-1/2x)+5

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

its supossed to be like this _+__________

OpenStudy (anonymous):

oh sorry, you said standard form. all you have to do is gat both x and y on the same side, so from y=(-1/2x)+5 just add (-1/2x) to both sides and your answer should be 1/2x+y=5

OpenStudy (anonymous):

so..5+1/2=y

OpenStudy (anonymous):

lol what jolly

OpenStudy (anonymous):

in standard form for a line, your equation should look like Ax+By=C so it would be 1/2x+y=5

OpenStudy (anonymous):

haha, they just dont let you put in fractions... so

OpenStudy (anonymous):

what did u get for #12 jolly?

OpenStudy (anonymous):

but cant u put 1(division sign)2

OpenStudy (anonymous):

does it work?

OpenStudy (anonymous):

idunno .-.

OpenStudy (anonymous):

oh, then it may be (.5)x+y=5

OpenStudy (anonymous):

no "."'s either

OpenStudy (anonymous):

then you could multiply the whole equation by (2) (2)(1/2x+y=5) and you get x+2y=10

OpenStudy (anonymous):

could you elmininate the fraction and make it x+y = 10?

OpenStudy (anonymous):

eliminate*

OpenStudy (anonymous):

just kidding:) i'm wrong.

OpenStudy (anonymous):

since you are eliminating a fraction from a term, remember to multiply every term, so that you have (y)(2)